# LoRA-IR: Taming Low-Rank Experts for Efficient All-in-One Image Restoration

### v1

@@ 4 Experiments
Method Dehazing Deraining Denoising on BSD68 dataset Average
on SOTS on Rain100L $\sigma=15$ $\sigma=25$ $\sigma=50$
@@
Method Rain Haze Snow Average
MPRNet [77] 21.50 21.73 20.74 21.32
NAFNet [4] 23.01 22.20 22.11 22.44
Uformer [65] 22.25 18.81 20.94 20.67
Restormer [78] 23.67 22.90 22.51 22.86
GRL [26] 23.75 22.88 22.59 23.07
AirNet [22] 22.52 21.56 21.44 21.84
TUM [6] 23.22 22.38 22.25 22.62
Transweather [60] 22.21 22.55 21.79 22.18
WGWS [95] 23.80 22.78 22.72 23.10
LDR [73] 24.42 23.11 23.12 23.55
LoRA-IR (Ours) 25.22 24.39 23.31 24.31
@@ 4.1 Experimental Setup
Settings. To comprehensively evaluate our method, we conduct experiments in five different settings following previous works: (I) 4-task adverse weather removal [56], including desnowing, deraining, dehazing, and raindrop removal; (II) 3-task real-world adverse weather removal [73], including deraining, dehazing, and desnowing; (III) 3-task image restoration [22], including deraining, dehazing, and denoising; (IV) 5-task image restoration [89], including deraining, low-light enhancement, desnowing, dehazing, and deblurring; (V) 10-task image restoration [36], including deblurring, dehazing, JPEG artifact removal, low-light enhancement, denoising, raindrop removal, deraining, shadow removal, desnowing, and inpainting. For each setting, we train a unified model to handle multiple types of degradations.
Datasets and Metrics. For Setting I, we use the AllWeather [60, 56] dataset to evaluate our method. For Setting II, we use the WeatherStream [80] dataset to evaluate the model’s performance in real-world scenarios. For Setting III, we use RESIDE [21] for dehazing, WED [39] and BSD [41] for denoising, Rain100L [74] for deraining. For Setting IV, we use a merged dataset [78, 89] for deraining, LOL [68], DCIE [19], MEF [38], and NPE [62] for low-light enhancement, Snow100K [34] for desnowing, RESIDE [21] for dehazing, GoPro [44], HIDE [55], RealBlur [52] for deblurring. For Setting V, we use the same dataset as [36]. We provide a detailed introduction to the datasets used for training and testing in Appendix.
@@
Implementation Details. For the training of DG-Router, we use the Adam optimizer with a batch-size of $64\times n$, where $n$ is the number of tasks. The whole training takes 20 minutes with a fixed learning rate of $2e^{-4}$ using 8 NVIDIA A100 GPUs. Our LoRA-IR follows a two-stage training process, namely pre-training and fine-tuning. For both stages, we use the AdamW optimizer with a batch-size of 64. Following [36, 89], the training patch size is set to $256$ to ensure fair comparisons. Random cropping, flipping, and rotation are used as data augmentation techniques. For the image restoration network structure, all basic blocks in Fig. 4 are the simple convolutional NAFBlocks [4], forming a simple all-in-one CNN baseline. More specific details for different settings provided in Appendix.
4.2 Comparison with State-of-the-Art Methods
@@
Method Deraining $(5sets)$ Enhancement Desnowing $(2sets)$ Dehazing Deblurring
@@
SwinIR [27] - - 17.81 0.723 - - 21.50 0.891 24.52 0.773
MIRNetV2 [79] - - 24.74 0.851 - - 24.03 0.927 26.30 0.799
IR-SDE [35] - - 20.45 0.787 - - - - 30.70 0.901
WeatherDiff [46] - - - - 33.51 0.939 - - - -
RDDM [30] 30.74 0.903 23.22 0.899 32.55 0.927 30.78 0.953 29.53 0.876
@@
Restormer [78] 27.10 0.843 17.63 0.542 28.61 0.876 22.79 0.706 26.36 0.814
AirNet [22] 24.87 0.773 14.83 0.767 27.63 0.860 25.47 0.923 26.92 0.811
Painter [64] 29.49 0.868 22.40 0.872 - - - - - -
IDR [81] - - 21.34 0.826 - - 25.24 0.943 27.87 0.846
ProRes [37] 30.67 0.891 22.73 0.877 - - - - 27.53 0.851
PromptIR [48] 29.56 0.888 22.89 0.847 31.98 0.924 32.02 0.952 27.21 0.817
DACLIP-UIR [36] 28.96 0.853 24.17 0.882 30.80 0.888 31.39 0.983 25.39 0.805
DiffUIR-L [89] 31.03 0.904 25.12 0.907 32.65 0.927 32.94 0.956 29.17 0.864
LoRA-IR (Ours) 32.35 0.924 26.42 0.926 34.16 0.941 35.74 0.986 30.09 0.914
@@
LoRA-IR (Ours) 28.68 0.876 0.279 17.02 0.700 0.600
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Low-light&Blurry Input DACLIP-UIR [36] DiffUIR [89] LoRA-IR (Ours) GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Rainy&Hazy Input DACLIP-UIR [36] DiffUIR [89] LoRA-IR (Ours) GT
@@
Setting I. Tab. 1 shows the comparison results of our method with task-specific methods and all-in-one methods. Compared to SOTA methods like MPerceiver [2] and Histoformer [56], our approach shows significant improvements across all benchmarks and metrics. Visual comparisons are provided in the Appendix.
Setting II. To further demonstrate the effectiveness of our method in mitigating real-world adverse weather conditions, we evaluate its performance on the WeatherStream [80] dataset. Tab. 3 presents the quantitative comparison results of PSNR with SOTA general IR as well as all-in-one IR methods. Compared to the SOTA method LDR [73], our method achieves an average PSNR improvement of 0.76 dB across the three tasks. Visual comparisons are provided in the Appendix.
Setting III. Tab. 2 presents the quantitative comparison results for 3-task image restoration. Our method surpasses PromptIR [48] by 1.38 dB in PSNR and 0.007 in SSIM on the Rain100L dataset, with an average improvement of 0.36 dB across the three tasks. Visual comparisons are provided in the Appendix.
Setting IV. Tab. 4 presents the quantitative comparison results of our method against SOTA task-specific methods and all-in-one methods across five tasks. It shows that our method outperforms the compared all-in-one methods across all tasks and even surpasses task-specific methods in the majority of tasks. For example, compared to the recent SOTA method DiffUIR [89], LoRA-IR brings a PSNR improvement ranging from 0.92 to 2.8 dB across various tasks.
(1) Generalization on Training-seen Tasks: We directly test the trained all-in-one model on real-world benchmarks that were not seen during training. As shown in Tab. 5, our method achieves the best PSNR and SSIM metrics in deblurring tasks. As discussed in  [76, 69], diffusion-based IR methods typically have an advantage in no-reference metrics like NIQE. However, our CNN-based model achieves comparable or even better performance on no-reference metrics compared to two SOTA diffusion-based methods, DACLIP-UIR [36] and DiffUIR [89]. Notably, our method shows approximately a 100-point improvement in LOE performance over DiffUIR in enhancement. Fig. 5 also shows that our method achieves more pleasing visual results.
@@
Mixed-degradataion Removal. Considering that images in real-world scenarios may not contain only a single type of degradation, we further evaluated different all-in-one methods on mixed-degradation benchmarks. Our experiments include three mixed-degradation benchmarks: rain&haze [23], low-light&blur [91], and blur&JPEG [45]. Tab. 9 shows that our method has a significant advantage in handling challenging mixed-degradation scenarios. We provide visual results in Fig. 6, showcasing the effectiveness of our method in handling mixed degradations.
@@
Snow [34] Rain [23] Raindrop [49] Mixed1[91] Mixed2[45]
@@ 5 Conclusion
This paper introduces LoRA-IR, a flexible framework that dynamically leverages compact low-rank experts to facilitate efficient all-in-one image restoration. We propose a CLIP-based degradation-guided router (DG-Router) to extract robust degradation representations, requiring minimal training parameters and time. With the valuable guidance of the DG-Router, LoRA-IR dynamically integrates different low-rank experts, enhancing architectural flexibility while preserving computational efficiency. Across 14 image restoration tasks and 29 benchmarks, LoRA-IR demonstrates its state-of-the-art performance and strong generalizability.

### v2

@@ 4 Experiments
Method Dehazing Deraining Denoising on BSD68 [41] Average
on SOTS [21] on Rain100L [74] $\sigma=15$ $\sigma=25$ $\sigma=50$
@@
Method Venue Rain Haze Snow Average
MPRNet [77] CVPR’21 21.50 21.73 20.74 21.32
NAFNet [4] ECCV’22 23.01 22.20 22.11 22.44
Uformer [65] CVPR’22 22.25 18.81 20.94 20.67
Restormer [78] CVPR’22 23.67 22.90 22.51 22.86
GRL [26] CVPR’23 23.75 22.88 22.59 23.07
AirNet [22] CVPR’22 22.52 21.56 21.44 21.84
TUM [6] CVPR’22 23.22 22.38 22.25 22.62
Transweather [60] CVPR’22 22.21 22.55 21.79 22.18
WGWS [95] CVPR’23 23.80 22.78 22.72 23.10
LDR [73] CVPR’24 24.42 23.11 23.12 23.55
LoRA-IR - 25.22 24.39 23.31 24.31
@@ 4.1 Experimental Setup
Settings. To comprehensively evaluate our method, we conduct experiments in five different settings following previous works: (I) 4-task adverse weather removal [56], including desnowing, deraining, dehazing, and raindrop removal; (II) 3-task real-world adverse weather removal [73], including deraining, dehazing, and desnowing; (III) 3-task image restoration [22], including deraining, dehazing, and denoising; (IV) 5-task image restoration [89], including deraining, low-light enhancement, desnowing, dehazing, and deblurring; (V) 10-task image restoration [36], including deblurring, dehazing, JPEG artifact removal, low-light enhancement, denoising, raindrop removal, deraining, shadow removal, desnowing, and inpainting. For each setting, we train a single model to handle multiple types of degradation.
Datasets and Metrics. For Setting I, we use the AllWeather [60, 56] dataset to evaluate our method. For Setting II, we use the WeatherStream [80] dataset to evaluate the model’s performance in real-world scenarios. For Setting III, we use RESIDE [21] for dehazing, WED [39] and BSD [41] for denoising, Rain100L [74] for deraining. For Setting IV, we use a merged dataset [78, 89] for deraining, LOL [68], DCIE [19], MEF [38], and NPE [62] for low-light enhancement, Snow100K [34] for desnowing, RESIDE [21] for dehazing, GoPro [44], HIDE [55], RealBlur [52] for deblurring. For Setting V, we use the same dataset as [36]. Due to space limitations, further information on the training dataset, training protocols, and additional visual results are provided in the Appendix.
@@
Implementation Details. For the training of DG-Router, we use the Adam optimizer with a batch-size of $64\times n$, where $n$ is the number of tasks. The whole training takes 20 minutes with a fixed learning rate of $2e^{-4}$ using 8 A100 GPUs. Our LoRA-IR follows a two-stage training process, namely pre-training and fine-tuning. For both stages, we use the AdamW optimizer with a batch-size of 64. Following [36, 89], the training patch size is set to 256 to ensure fair comparisons. Random cropping, flipping, and rotation are used as data augmentation techniques. For the pre-training stage, we use an initial learning rate of $1e^{-3}$, which is updated using the cosine annealing scheduler after 200000 iterations. The minimal learning rate is set to $1e^{-5}$. For fine-tuning, we use an initial learning rate of $1e^{-4}$, which decreases to $1e^{-5}$ after 100000 iterations. For the image restoration network structure, all basic blocks in Fig. 4 are the simple convolutional NAFBlocks [4], forming a simple all-in-one CNN baseline. More specific details for different settings are provided in the Appendix.
4.2 Comparison with State-of-the-Arts
@@
Method Venue Deraining $(5sets)$ Enhancement Desnowing $(2sets)$ Dehazing Deblurring
@@
SwinIR [27] ICCVW’21 - - 17.81 0.723 - - 21.50 0.891 24.52 0.773
MIRNetV2 [79] TPAMI’22 - - 24.74 0.851 - - 24.03 0.927 26.30 0.799
IR-SDE [35] ICML’23 - - 20.45 0.787 - - - - 30.70 0.901
WeatherDiff [46] TPAMI’23 - - - - 33.51 0.939 - - - -
RDDM [30] CVPR’24 30.74 0.903 23.22 0.899 32.55 0.927 30.78 0.953 29.53 0.876
@@
Restormer [78] CVPR’22 27.10 0.843 17.63 0.542 28.61 0.876 22.79 0.706 26.36 0.814
AirNet [22] CVPR’22 24.87 0.773 14.83 0.767 27.63 0.860 25.47 0.923 26.92 0.811
Painter [64] CVPR’23 29.49 0.868 22.40 0.872 - - - - - -
IDR [81] CVPR’23 - - 21.34 0.826 - - 25.24 0.943 27.87 0.846
ProRes [37] arXiv’23 30.67 0.891 22.73 0.877 - - - - 27.53 0.851
PromptIR [48] NeurIPS’23 29.56 0.888 22.89 0.847 31.98 0.924 32.02 0.952 27.21 0.817
DACLIP-UIR [36] ICLR’24 28.96 0.853 24.17 0.882 30.80 0.888 31.39 0.983 25.39 0.805
DiffUIR-L [89] CVPR’24 31.03 0.904 25.12 0.907 32.65 0.927 32.94 0.956 29.17 0.864
LoRA-IR - 32.35 0.924 26.42 0.926 34.16 0.941 35.74 0.986 32.05 0.927
@@
LoRA-IR 28.68 0.876 0.279 17.02 0.700 0.600
Setting I. Tab. 1 shows the comparison results with task-specific methods and all-in-one methods. Compared to SOTA methods like MPerceiver [2] and Histoformer [56], our approach shows significant improvements across all benchmarks and metrics.
Setting II. To further demonstrate the effectiveness of our method in mitigating real-world adverse weather conditions, we evaluate its performance on the WeatherStream [80] dataset. Tab. 3 presents the quantitative comparison results of PSNR with SOTA general IR as well as all-in-one IR methods. Compared to the SOTA method LDR [73], our method achieves an average PSNR improvement of 0.76 dB across the three tasks.
Setting III. Tab. 2 presents the quantitative comparison results for 3-task image restoration. Our method surpasses PromptIR [48] by 1.38 dB in PSNR on the Rain100L dataset, with an average improvement of 0.36 dB across the three tasks.
Setting IV. Tab. 4 presents the quantitative comparison results of our method against SOTA task-specific methods and all-in-one methods across five tasks. It shows that our method outperforms the compared all-in-one and task-specific methods across all tasks. For example, compared to the recent SOTA all-in-one method DiffUIR [89], LoRA-IR brings a PSNR improvement ranging from 0.92 dB to 2.87 dB across various tasks.
(1) Generalization on Training-seen Tasks: We directly test the trained all-in-one model on real-world benchmarks that were not seen during training. As shown in Tab. 5, our method achieves the best PSNR and SSIM metrics for deblurring. As discussed in  [76, 69], diffusion-based IR methods typically have an advantage in no-reference metrics like NIQE. However, our CNN-based model achieves comparable or even better performance on no-reference metrics compared to two SOTA diffusion-based methods, DACLIP-UIR [36] and DiffUIR [89]. Notably, our method shows approximately a 100-point improvement in LOE performance over DiffUIR in enhancement. Fig. 5 also shows that our method achieves more pleasing visual results.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Low-light&Blurry Input PromptIR [48] DACLIP-UIR [36] DiffUIR [89] LoRA-IR (Ours) GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Rainy&Hazy Input PromptIR [48] DACLIP-UIR [36] DiffUIR [89] LoRA-IR (Ours) GT
@@
Mixed-degradataion Removal. Considering that images in real-world scenarios may not contain only a single type of degradation, we further evaluate different all-in-one methods on mixed-degradation benchmarks. Our experiments include three mixed-degradation benchmarks: rain&haze [23], low-light&blur [91], and blur&JPEG [45]. Tab. 9 shows that our method has a significant advantage in handling challenging mixed-degradation scenarios. We provide visual results in Fig. 6, showcasing the effectiveness of our method in handling mixed degradations.
@@
LoRA-IR Snow [34] Rain [23] Raindrop [49] Mixed1[91] Mixed2[45]
@@ 5 Conclusion
This paper introduces LoRA-IR, a flexible framework that dynamically leverages compact low-rank experts to facilitate efficient all-in-one image restoration. We propose a CLIP-based Degradation-Guided Router (DG-Router) to extract robust degradation representations, requiring minimal training parameters and time. With the valuable guidance of the DG-Router, LoRA-IR dynamically integrates different low-rank experts, enhancing architectural adaptability while preserving computational efficiency. Across 14 image restoration tasks and 29 benchmarks, LoRA-IR demonstrates its state-of-the-art performance and strong generalizability.